ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

9. Ordinary Differential Equations and Dynamical Systems

1991 MS Classification: 34D08, 34C35, 58F10 Rangarajan, Govindan, Indian Institute of Science, India A New Method for Computing Lyapunov Exponents We present a new method for the computation of Lyapunov exponents utilizing representations of orthogonal matrices applied to decompositions of M or MM̃ where M is the tangent map. This method uses a minimal set of variables, does not require renormalization or reorthogonalization, can be used to efficiently compute partial Lyapunov spectra, and does not break down when the Lyapunov spectrum is degenerate. Since our method is based on exact differential equations for the Lyapunov exponents, global invariances of the Lyapunov spectrum can be preserved.